OpenStudy (anonymous):

At a certain temperature, molecules of bromine gas travel at 0.300Km/s. What is the speed of CH2O gas molecules under the same condition?

OpenStudy (aaronq):

use grahams law of effusion \(\dfrac{v_1}{v_2}=\sqrt{\dfrac{M_2}{M_1}}\) M=molar mass